1Kg of real gas is heated from 18ºC to 93 ºC. Assuming Cv=1.466KJ/KgK. Then internal energy is ……

1
Expert's answer
2021-11-11T11:59:03-0500

For real gas "U=\nu C_VT-a\nu^2/V" . So, if "V=const" then the change of


the internal energy of real gas is


"\Delta U= \nu C_V\Delta T=\frac{m}{M} c_VM\Delta T=m c_V\Delta T="


"=1\cdot 1466\cdot(93°-18°)=109950\ (J)" . Answer